OpenStudy (migitmack):

Solve 2x^2 + 3x + 6 = 0. Round solutions to the nearest hundredth.

Parth (parthkohli):

Have you heard of the quadratic formula?

OpenStudy (migitmack):

yyeah

OpenStudy (unklerhaukus):

\[ax^2+bx+c=0\qquad\Longrightarrow\qquad x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

Parth (parthkohli):

\[ \color{Black}{\Rightarrow 2x^2 + 3x + 6 = 0 \Longrightarrow {x = {-3 \pm \sqrt{3^2 - 4(2)(6) } \over 2(2)}} }\]